# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

### v1

@@ 4 Related Work
The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as (Lampert et al., 2009) and (Kumar et al., 2009). A first systematic study of CBMs and their variations was delivered by (Koh et al., 2020) and (Stammer et al., 2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output. Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$. Since the initial works on CBMs they have found utilization in several applications. (Antognini & Faltings, 2021) e.g. apply CBMs to text sentiment classification. (Bahadori & Heckerman, 2020) propose a bias-removing method to tackle the problem of confounding concept correlations. (Chauhan et al., 2022) on the other hand query a domain expert for concept adaptation at prediction time. However, all of these works consider only single object concept representations and mainly focus on supervised learning settings.
@@
In fact, CBMs have found their way into RL only to a limited extent. (Zabounidis et al., 2023) and (Grupen et al., 2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance. (Zabounidis et al., 2023) further report better training stability and reduced sample complexity. Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part. SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations. Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, (Guan et al., 2023) allow the user to design a reward function based on “behavioral attributes”, i.e. higher-level properties that occur not in a single state but over a period of time. Lastly, SCoBots not only separate state representation learning from policy search, as done by (Cuccu et al., 2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
@@
Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022). For example, the true qualities of a good explanation are unclear and there is a lack of comparative explainability evaluation benchmarks. (Milani et al., 2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior. Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods. According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining). Moreover, (Dazeley et al., 2021) suggest different “levels of explanation” based on the agent’s intentions. SCoBots especially facilitate the generation of zero-order explanations, i.e., reactive explanations that explain the agent’s reaction of an immediately perceived input. Overall, SCoBots transform the raw state $s$ into a symbolic state $s^{\prime}=\mu(\omega(s))$ in line with the proposal of Kambhampati et al.. d’Avila Garcez et al. do an analogous symbolic extraction but resulting in a set of symbolic “sub-states” $s_{k}^{\prime}$ instead of $s^{\prime}$: Each $s_{k}^{\prime}$ represents the relative position of the agent to one interactable object. This way, however, solving one MDP results in solving a set of partially observable MDPs which is avoided in the case of SCoBots.
@@
The misalignment problem is related to shortcut learning, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023) to prototype-based models (Bontempelli et al., 2023) to reinforcement learning (RL) approaches (di Langosco et al., 2022). A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022). The authors (Nahian et al., 2021) align the agent to ethical values by introducing a second reward signal. In comparison, SCoBots aid to resolve RL specific learning issues such as the misalignment problem through inherent interpretability.
@@ 5 Conclusion
In this work, we have shown the benefits of concept-based models in RL tasks, specifically for identifying issues such as goal misalignment. Among other things our proposed Successive Concept Bottleneck agents specifically integrate automatically generated relational concept representations into their decision processes. With this we could identify previously unknown misalignment problems of deep RL agents in games as simple as Pong. Due to their inspectable and human-understandable nature of our proposed SCoBot agents these issues could not only be identified, but easily revised with minimal additional feedback. Our work thus represents an important step in developing aligned RL agents, i.e. agents that are not just aligned with the underlying task goals, but also with human user’s understanding and knowledge. Achieving this is particularly valuable for applying RL agents in real-world settings where ethical and safety considerations are paramount.
Avenues for future research are incorporating options, which represent higher-level actions as described in (Bacon et al., 2017), in order to make our (i)SCoBots more closely resemble human-like decision-making processes. One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023). Additionally, we are considering the use of shallower decision trees, as proposed by (Broelemann & Kasneci, 2019), to enhance interpretability further in our RL agents. In our experiments we assume the initial object concept representations have been extracted. An interesting research question is further in how far the task reward signal can aid in learning such object-centric representations in the first place e.g. based on the MOC framework of Delfosse et al..
@@ Appendix A Appendix
As mentioned in the main body, the appendix contains additional materials and supporting information for the following aspects: further information on the fact that Atari is the most common set of games (A.1), details on the reward sparsity in Pong (A.2), detailed numerical results (A.3), learning curves of our RL agents (A.4), the hyperparmeters used in this work (A.6), formal definitions on the SCoBots policies (A.7) and further details on the misalignment problem in Pong.
@@ A.1 Atari games are most common set of environments
A.2 The reward sparsity of Pong.
In the game Pong, the agent has a low chance of discovering positive reward. Its vertical position may vary between $34$ and $190$. Its paddle height is $16$, but it needs to shoot from the side of the paddle (the $3$ pixel of each border) to get a spiky shot that is likely to go beyond the opponent. If the balls arrives not close to the top or bottom borders, the enemy will still catch it, which has the probability of $\sim 22\%$ in our experiments. The probability of getting a successful shot is thus of $\sim 6/156*0.78=3\%$. The average number of corresponding tryouts for the agent to get rewarded is $\sum_{n=1}^{\infty}\left(n\times 0.03\times 0.97^{n-1}\right)=33.3$. The agent usually need $60$ steps (in average) from the initialization of the point to the reward attribution, hence, it will need $\sim 2200$ steps to be rewarded. This is consistent with the experiment depicted in Figure 6. In this figure, a random agent (original) needs in average $2230$ steps to observe reward.
By providing the agent with a punishment proportional to the vertical distance between its paddle and the ball, we were able to obtain a winning shot every $\sim\!820$, multiplying by $2.7$ their occurrences.
A.3 Detailed numerical results
For completeness, we here provide the numerical results of the performances obtained by each agent type. For fair comparison, we reimplemeneted the deep PPO agents, and used the default hyperparameters for Atari from stable-baselines3 for the deep RL agents. A detailed overview of hyperparameters and their corresponding values can be found in A.6.
@@ A.3 Detailed numerical results
Game SCoBots SCoBots (guided) Neural(v5) Original reproduced (v4) Original Random Human
@@
Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs. Deep refers to the agent that use a CNN instead of our ICB layers. We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)). Our agents have been trained with only 2020M frames.
@@ A.4 Learning curves
A.5 Agent reasoning through the decision trees.
@@
A.6 Hyperparameters and experimental details
@@
Table 2: PPO Hyperparameter Values. $\alpha$ linearly decreases from 11 to 0 over the course of training.
@@
The Atari environment version used in gymnasium is NoFrameskip-v4 for agents reproducing the reported PPO results, and v5 for SCoBots and neural. v5 defines a deterministic skipping of 5 frames per action taken and sets the probability to repeat the last action taken to 0.25. This is aligned with recommended best practices by Machado et al. (2018). The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation222github.com/openai/baselines. This includes frame skipping of 4 and reward clipping. Frame stacking is not used. SCoBots are not trained on a clipped reward signal. For comparability, the neural agents we compare SCoBots to, are not as well. A list of all hyperparameter values used is provided in Table 2.
A.6.1 The properties and features used for SCoBots
@@
color $color(o_{1})$ the CSS2.1333https://www.w3.org/TR/CSS2/ color category (e.g. Red)
@@
A.7 SCoBots policies: formal definitions
@@
A.8 Pong misalignment problem.
The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019). Here, an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time. When the coin is repositioned at test time di Langosco et al. observed that trained deep RL agents avoid the coin and simply target the end of the level, indicating that the agents have in fact learned to follow a simpler side-goal rather than the underlying strategy.
To find the correlation between the enemy’s and the ball’s y positions, we let an random agent play the game for $100000$ frames and collect the positions of the enemy and ball every $10$ frames. We then compute different correlation coefficients, and obtain 99.6% as Pearson coefficient, 96.4% as Kendall coefficient and 99.5% as Spearman coefficient.
Here are then importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2021).
@@ A.9 RL specific caveats
Sparse Reward: Some RL environments provide the agent with a rarely occurring reward signal, making it challenging to learn effectively. Such sparse reward environments are particularly present in the robotic field (Andrychowicz et al., 2017).
Ill-defined Objectives: Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018). Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common (Andrychowicz et al., 2017).
@@
Difficult Credit Assignment: When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021). This problem arises in games like chess, where reward is passed to the agent only when the game is over. The problem can be addressed by more instant informative feedback, given directly after successful actions or after a positive state is reached.
@@
This results in an agent learning to land close to the Flags (cf. Fig. 12)
To avoid this behavior, we add another signal, based on the speed of the agent:
@@
Misalignment: Reward shaping permits realigning the agent on the true task objectives, preventing undesirable behaviors resulting from misaligned optimization. Shortcut learning can be mitigated through the implementation of reward signals that necessitate the genuine objective’s consideration.

### v2

@@ 5 Related Work
The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as Lampert et al. (2009) and Kumar et al. (2009). A first systematic study of CBMs and their variations was delivered by Koh et al. (2020) and Stammer et al. (2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output. Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$. Since the initial works on CBMs they have found utilization in several applications. Antognini & Faltings (2021) e.g. apply CBMs to text sentiment classification. However, these works consider single object concept representations and focus on supervised learning.
@@
In fact, CBMs have found their way into RL only to a limited extent. Zabounidis et al. (2023) and Grupen et al. (2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance. Zabounidis et al. (2023) further report better training stability and reduced sample complexity. Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part. SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations. Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, Guan et al. (2023) allow the user to design a reward function based on higher-level properties that occur over a period of time. Lastly, SCoBots not only separate state representation learning from policy search, as done by Cuccu et al. (2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
@@
Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022). Milani et al. (2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior. Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods. According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining).
@@
The misalignment problem is an RL instantiation of the shortcut learning problem, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023), prototype-based models (Bontempelli et al., 2023) and RL approaches (di Langosco et al., 2022). A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022). Nahian et al. (2021) ethically align agents by introducing a second reward signal. In comparison, SCoBots aid to resolve RL specific issues such as the misalignment problem through inherent interpretability.
@@ 6 Conclusion
In this work, we have provided evidence for the benefits of concept-based models in RL tasks, specifically for identifying issues such as goal misalignment. Among other things our proposed Successive Concept Bottleneck agents integrate relational concepts into their decision processes. With this, we have exposed previously unknown misalignment problems of deep RL agents in a game as simple as Pong. SCoBot agents allowed us to revise this issue, as well as different RL specific caveats with minimal additional feedback. Our work thus represents an important step in developing aligned RL agents, i.e. agents that are not just aligned with the underlying task goals, but also with human user’s understanding and knowledge. Achieving this is particularly valuable for applying RL agents in real-world settings where ethical and safety considerations are paramount.
Avenues for future research are incorporating a high level action bottleneck (Bacon et al., 2017). One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023). Additionally, we are considering the use of shallower decision trees (Broelemann & Kasneci, 2019). An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023c) in the first place.
Impact statement
Our work aims at developing transparent RL agents, whose decision can be understood and revised to be aligned with the beliefs and values of a human user. We believe that such algorithms are critical to uncover and mitigate potential misalignments of AI systems. A malicious user can, however, utilize such approaches for aligning agents in a harmful way, thereby potentially leading to a negative impact on further users or society as a whole. Even so, the inspectable nature of transparent approaches will allow to identify such potentially harmful misuses, or hidden misalignment.
@@ Appendix A Appendix
As mentioned in the main body, the appendix contains additional materials and supporting information for the following aspects: further information on the fact that Atari is the most common set of games (A.1), details on the reward sparsity in Pong (A.8.3), detailed numerical results (A.2), learning curves of our RL agents (A.3), the hyperparmeters used in this work (A.5), formal definitions on the SCoBots policies (A.6) and further details on the misalignment problem in Pong.
@@ A.1 Atari games are most common set of environments
A.2 Detailed numerical results
For completeness, we here provide the numerical results of the performances obtained by each agent type. For fair comparison, we reimplemeneted the deep PPO agents, and used the default hyperparameters for Atari from stable-baselines3 for the deep RL agents. A detailed overview of hyperparameters and their corresponding values can be found in A.5.
@@ A.2 Detailed numerical results
Game
SCoBots
v5
SCoBots
guided-v5
PPO
ours-v5
PPO
ours-v4
PPO
Schulman et al.-v4
Random Human
@@
Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs. Neural refers to the agent that use a CNN instead of our ICB layers. We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)). Our agents have been trained with only 2020M frames.
@@ A.3 Learning curves
A.4 Agent reasoning through the decision trees
@@
A.5 Hyperparameters and experimental details
@@
Table 2: PPO Hyperparameter Values. $\alpha$ linearly decreases from $1$ to 0 over the course of training.
@@
The Atari environment version used in gymnasium is NoFrameskip-v4 for agents reproducing the reported PPO results, and v5 for SCoBots and neural. v5 defines a deterministic skipping of 5 frames per action taken and sets the probability to repeat the last action taken to 0.25. This is aligned with recommended best practices by Machado et al. (2018). The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation444github.com/openai/baselines. This includes frame skipping of 4 and reward clipping. Frame stacking is not used. SCoBots are not trained on a clipped reward signal. For comparability, the neural agents we compare SCoBots to, are not as well. A list of all hyperparameter values used is provided in Table 2.
A.5.1 The properties and features used for SCoBots
@@
color $color(o_{1})$ the CSS2.1555https://www.w3.org/TR/CSS2/ color category (e.g. Red)
@@
A.6 SCoBots policies: formal definitions
@@
A.7 Pong misalignment problem
The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019), in which an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time. When the coin is repositioned at test time di Langosco et al. observed that trained deep RL agents avoid the coin and simply target the end of the level, indicating that agents in fact learn to follow a simpler side-goal rather than the underlying strategy.
For the correlation between the enemy’s and the ball’s y positions, we let an random agent play the game for $100000$ frames and collect the positions of the enemy and ball every $10$ frames. We then compute different correlation coefficients, and obtain 99.6% as Pearson coefficient, 96.4% as Kendall coefficient and 99.5% as Spearman coefficient.
We also collected importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2024).
@@ A.8 RL specific caveats
A.8.1 Ill-defined Objectives
Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018). Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common (Andrychowicz et al., 2017).
@@ A.8.1 Ill-defined Objectives
A.8.2 Difficult Credit Assignment
When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021). This problem arises in games like chess, where reward is passed to the agent only when the game is over. The problem can be addressed by more instant informative feedback, given directly after successful actions or after a positive state is reached.
@@ A.8.2 Difficult Credit Assignment
A.8.3 The reward sparsity of Pong
Let us move on to the issue of sparse reward in the context of RL. In Pong, to score a point, the player needs to return the ball and have it go past the enemy. To do so, the player has to perform a spiky shot, obtained by touching the ball on one of its paddle’s sides (otherwise, the shot is flat and the enemy is easily catching it). Its vertical position may vary between $34$ and $190$. Its paddle height is $16$, but it needs to shoot from the side of the paddle (the $3$ pixel of each border) to get a spiky shot that is likely to go beyond the opponent. If the balls arrives not close to the top or bottom borders, the enemy will still catch it, which has the probability of $\sim 22\%$ in our experiments. The probability of getting a successful shot is thus of $\sim 6/156*0.78=3\%$. The average number of corresponding tryouts for the agent to get rewarded is $\sum_{n=1}^{\infty}\left(n\times 0.03\times 0.97^{n-1}\right)=33.3$. The agent usually need $60$ steps (in average) from the initialization of the point to the reward attribution, hence, it will need $\sim 2200$ steps to be rewarded. This is consistent with the experiment depicted in Figure 12. In this figure, a random agent (original) needs in average $2230$ steps to observe reward.
Providing a SCoBot agent with an additional reward that is inversely proportional to the distance between its paddle and the ball incentivizes the agent to keep a vertical position close to the ball’s one:
$R^{\text{exp}}=D(Player,Ball).y$ (11)
This extra reward signal lets a starting agent a winning shot every $\sim\!820$, multiplying by $2.7$ their occurrences. Thus, the reward signal in Pong is relatively sparse, but our example illustrates how the interpretable concepts allow to easily guide SCoBots, making up for the reward sparsity. Interestingly, this also giving a clearer incentive to the agent to concentrate on the ball’s position, and not on the enemy’s one. We observe that such agent do not rely on the enemy to master the game. Alternative techniques such as prioritized experience replay (Schaul et al., 2016) allow offline methods to learn in such environments, but providing a smoother reward signal is another elegant way to address sparsity. Note that sparser environments exists (such as robotics ones).
A.8.4 Misalignment
In our experimental evaluations, we used the concept based reward shaping to realign the agents on the true (target) task objectives, preventing undesirable behaviors resulting from misaligned optimization. Shortcut learning can be mitigated through the implementation of such reward signals that necessitate the genuine objective’s consideration.